OpenStudy (anonymous):

ok let me clarify.. help me refresh this to my mind.. (2x + 1)/(4x − 2) = (3x+5)/(6x-14) solve for x..

OpenStudy (anonymous):

yes you solve for X

OpenStudy (anonymous):

http://www.mathwarehouse.com/algebra/ scroll down and type your question there

OpenStudy (anonymous):

thank you very much.. this will help me.. thank you..

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wait what is the meaning of always true in some of the answers..?

OpenStudy (anonymous):

show me the question

OpenStudy (anonymous):

1/6x-1=4/(24x-4) and this is what i get when i input it.. always true..?

OpenStudy (anonymous):

does it say solve for x next to answer?

OpenStudy (anonymous):

at the top of the white box make sure it says algebra

OpenStudy (anonymous):

yep.. and still the final answer is always true.. but i forgot a another question with that equation.. it says that If there are infinitely many solutions, give the restrictions on x as a comma separated list.. there.. how will i deal with that..?

OpenStudy (anonymous):

or this question is somehow hard to understand.. hmm.. how will i clarify this.. hmm.. it says after this equation 1/6x-1=4/(24x-4) this follows If there are infinitely many solutions, give the restrictions on x as a comma separated list.. still confusing..? heheh